Find the missing term in harmonic sequence. Show the solution
___,1/2,2/5,1/3,___

Found 2 solutions by robertb, greenestamps:
Answer by robertb(5830)


If the numbers a, 1/2, 2/5, 1/3, b are in harmonic progression, then their reciprocals are in arithmetic progression, i.e.,

1/a, 2, 5/2, 3, 1/b

are in AP.

It is quite clear that 1/a = 3/2 and 1/b = 7/2. (The common difference is 1/2.)

===> a = 2/3 and b = 2/7.

Answer by greenestamps(13332)



The response from tutor @robertb shows one way of solving a problem like this involving harmonic sequences.

Here is a different way that I personally prefer.

In a harmonic sequence, the terms can be written with a common NUMERATOR, and with DENOMINATORS that form an arithmetic sequence.

So given the harmonic sequence

____, 1/2, 2/5, 1/3, ____

rewrite all the given terms with the same numerator:

____, 2/4, 2/5, 2/6, ____

Then the pattern is clear:

2/3, 2/4, 2/5, 2/6, 2/7

ANSWER: The missing numbers are 2/3 and 2/7